when looking on a graph and one dot is on (0,1) the other dot is on (4,8) what is the slope of the line?
(the line starts at the bottom left and goes toward the top right)
(Type an integer or simplified fraction. Type N is the slope is undefined)

Answer by tutor_paul(427)

When given 2 points on a line and asked to find the slope, use the following formula:
slope=(y2-y1)/(x2-x1)
In this case:
slope=(8-1)/(4-0)
slope = 7/4
================
Good Luck,
tutor_paul@yahoo.com